CTA deployment meeting

Europe/Zurich
600/R-001 (CERN)

600/R-001

CERN

4
Show room on map
Michael Davis (CERN)
    • 15:00 15:10
      EOSCTA Post-Migration Issues 10m

      DUNE

      • See #293 Migrate DUNE from CASTOR to CTA
      • Reclaim NA62 tapes I70019 and L75667
      • No significant activity expected before Data Challenge 3Q2021

      LHCb

      • #306 DAQ integration and testing, including FTS Archive Monitoring
      • LHCb reprocessing campaign (begins in May)

      NA62

      • Protect DAQ user from deleting files written by DAQ. #295 Fix ACLs for NA62 on CTA PUBLIC, #297 Testing of p and !d eos permissions, #304 Testing of !d ACL.
      • #291 T2 transfers are working because the firewall was opened, but this is a temporary situation. Any WLCG site wishing to access CTA must be on LHCONE.
    • 15:10 15:20
      CTA PUBLIC Migration from CASTOR 10m

      TOTEM

      • Migration this week, see #277
      • TOTEM users (members of group zj) should be able to read, but who does the prepare?
      • Post-migration: 0.5 PB of TOTEM files in EOS need a tape copy

      n_TOF

      • Tue 20 April 14.00: meeting with n_TOF to agree migration schedule
      • One blocking issue: recall to spinner space immediately after archiving
      • #143 nTOF tests on EOSCTAPUBLIC PPS , #161 Spinners space on EOSCTA PPS

      CLIC/ILC

      • #310 CLIC/ILC tests on EOSCTAPUBLIC PPS
      • Status update: progress of tests
      • Migration 1st week in June (provisional)

      COMPASS

      • #69 COMPASS tests on EOSCTAPUBLIC PPS
      • COMPASS snaphot imported to eosctapps, instance needs spinners added to start tests

      AMS

      • #82 AMS tests on EOSCTAPUBLIC PPS
      • Status update on tests
      • He says after these tests he can proceed with integration work on his side and prepare for the migration.
      • On our side, 125 tapes to repack, but 118 of these overlap with the CLIC/ILC tapes.

      NA61/SHINE

      • NA61/SHINE: setting up their DAQ system + storage. No schedule for testing yet.
      • Decouple online and offline workflows, migrate offline to CTA.
      • Start offline testing.

      Backup

      • #307 Migrate CASTOR Backup use cases to CTA
      • Status of encryption for CASTOR backup use cases
      • /afsmigration #282.

      Other Use Cases

      • r_public_misc: covers most of the rest of the small experiment use cases. Needs repacking to separate from r_public_user.
      • LEP tapepools: postponed until more files have been repacked.
    • 15:20 15:30
      CTA PUBLIC New Use Cases 10m
      • Vlado's instructions have been published as KB0007266 Archiving data in the CERN Tape Archive (CTA)
      • #271 Spacal
      • #274 FASER experiment on CTA
      • AWAKE RQF1735636 - Archive 13 TB of AWAKE Run-1 data in /eos/experiment/awake
      • AMBER (replacement for COMPASS at NA66)
    • 15:30 15:40
      CTA Software 10m
      • #312 Postgres performance tests
      • Upgrade to PostgreSQL 12
      • Performance monitoring dashboard
      • Review of tickets tagged Tape Server
    • 15:40 15:45
      AOB 5m
      • Michael will contact Fermi to propose a technical meeting in the near future